# BlackboxBench: A Comprehensive Benchmark of Black-box Adversarial Attacks

### v1

@@ E-A Description of static defenses
E-B Description of dynamic defenses
@@
Appendix F Description of analytical tools
In this section, we provide implementation details of our analytical tools mentioned in Sec. III-C.
@@ Appendix F Description of analytical tools
FullGrad is a variant of a saliency map that not only captures the importance of each pixel but also provides that of groups of pixels, which correspond to structures. Consider a neural network function $f:\mathbb{R}^{D}\rightarrow\mathbb{R}$ with inputs $\mathbf{x}\in\mathbb{R}^{D}$. Let $c_{l}$ be the channel of layer $l$ in a neural network. Let $\psi\left(\cdot\right)$ be the post-processing operations. The FullGrad saliency map is given by
$S_{f}(\mathbf{x})=\psi\left(\nabla_{\mathbf{x}}f(\mathbf{x})\odot\mathbf{x}\right)+\sum_{l\in L}\sum_{c\in c_{l}}\psi\left(f^{b}(\mathbf{x})_{c}\right),$ (8)
where $\nabla_{\mathbf{x}}f(\mathbf{x})$ is the input-gradient, $f^{b}(\mathbf{x})$, short-hand notation for $\nabla_{b}f(\mathbf{x},\mathbf{b})\odot\mathbf{b}$, is the bias-gradient. Together, they constitute full-gradients.
@@
$F(\tilde{\boldsymbol{x}})=f(\boldsymbol{x})=f\left(\mathcal{F}^{-1}(\tilde{\boldsymbol{x}})\right).$ (9)
@@
$S_{F}(\tilde{\boldsymbol{x}})$ $=\frac{\partial F(\tilde{\boldsymbol{x}})}{\partial\tilde{\boldsymbol{x}}(u,v,c)}$ (10)
@@
Adversarial divergence measurement tool When harnessing adversarial transferability from transfer-based attacks, the performance is notably destroyed by the misalignment between surrogate model $f_{s}$ with parameter $\theta_{s}$ and target models $f_{t}$ with parameter $\theta_{t}$. To properly measure this misalignment, we consider conditional adversarial distribution (CAD), denoted as $\mathcal{P}(\boldsymbol{\delta}_{\boldsymbol{x}}|\boldsymbol{x},\theta)$. As CAD models the distribution of perturbations $\boldsymbol{\delta}_{\boldsymbol{x}}$ conditioned on clean examples $\boldsymbol{x}$ for each model with parameter $\theta$, minimal divergence between CADs ensures a high transferability between models. The rationle behind is that if $\mathcal{P}(\boldsymbol{\delta}_{\boldsymbol{x}}|\boldsymbol{x},\theta_{s})$ and $\mathcal{P}(\boldsymbol{\delta}_{\boldsymbol{x}}|\boldsymbol{x},\theta_{t})$ yield similar probabilities for perturbation $\boldsymbol{\delta}_{\boldsymbol{x}}$, then the perturbation craft by $f_{s}$ can effectively substitute the authentic perturbation crafted by $f_{t}$, thus misleading $f_{t}$. Building on this premise, BlackobxBench quantifies this divergence between $\mathcal{P}(\boldsymbol{\delta}_{\boldsymbol{x}}|\boldsymbol{x},\theta_{t})$ and $\mathcal{P}(\boldsymbol{\delta}_{\boldsymbol{x}}|\boldsymbol{x},\theta_{s})$ as Adversarial Divergence (AD):
@@
$\text{AD}=1-\frac{1}{\log_{2}\left[\mathbb{E}_{\mathcal{P}(\boldsymbol{\delta}_{\boldsymbol{x}}\mid\boldsymbol{x},\theta_{s})}\left[\log\frac{\mathcal{P}(\boldsymbol{\boldsymbol{\delta}_{\boldsymbol{x}}}\mid\boldsymbol{x},\theta_{s})}{\mathcal{P}(\boldsymbol{\boldsymbol{\delta}_{\boldsymbol{x}}}\mid\boldsymbol{x},\theta_{t})}\right]\right]},$ (11)
where $\left[\mathbb{E}_{\mathcal{P}(\boldsymbol{\delta}_{\boldsymbol{x}}\mid\boldsymbol{x},\theta_{s})}\left[\log\frac{\mathcal{P}(\boldsymbol{\boldsymbol{\delta}_{\boldsymbol{x}}}\mid\boldsymbol{x},\theta_{s})}{\mathcal{P}(\boldsymbol{\boldsymbol{\delta}_{\boldsymbol{x}}}\mid\boldsymbol{x},\theta_{t})}\right]\right]$ denotes the KL divergence between $\mathcal{P}(\boldsymbol{\delta}_{\boldsymbol{x}}|\boldsymbol{x},\theta_{t})$ and $\mathcal{P}(\boldsymbol{\delta}_{\boldsymbol{x}}|\boldsymbol{x},\theta_{s})$. Adversarial Divergence normalizes the KL divergence to be bounded within [0, 1].
Following [18], we learned the conditional generative flow model (c-Glow) for each classifier $f$ to approximate its CAD. The training details can be found in the supplementary materials in [18]. With two CADs, we compute the KL divergence between them as follows:
@@
    $\log\frac{\mathcal{P}(\boldsymbol{\delta}_{\boldsymbol{x}}^{(i)}\mid\boldsymbol{x},\theta_{s})}{\mathcal{P}(\boldsymbol{\delta}_{\boldsymbol{x}}^{(i)}\mid\boldsymbol{x},\theta_{t})}.$ (12)
@@
Decision boundary visualization tool Decision Surface[120] measures the network’s loss variations concerning the input perturbations and contains the explicit decision boundary, naturally aligning with our analytical requirements of measuring the distance between the benign images and decision boundary.
@@
$V(i,j,\alpha,\beta)=F(o+i\cdot\alpha+j\cdot\beta).$ (14)
For our purpose of assessing the network’s loss variations w.r.t perturbations, $\alpha$ represents an attack-specific perturbation direction, $\beta$ stands for a random direction vector, and $o$ denotes the benign image. The variables $i$ and $j$ quantify the degree of deviation from point $o$ along the $\alpha$ and $\beta$ directions, respectively.
@@
As shown in Fig. 10, with the loss for clean images serving as the central point, the decision surface show the loss variation w.r.t the perturbation density (shown in x-axis) and the noise density along (shown in y-axis). Evidently, it requires a shorter distance to traverse the decision boundary in the adversarial direction compared to the random direction. This observation indicates that adversarial attacks can identify directions toward the decision boundary within the immediate vicinity of clean images.
@@
Appendix G Additional evaluation results
G-A Intra/Cross-family transferability
In Sec. IV-E, we explore the transferability between models from four families employing two mechanisms. With the help of DI-FGSM[25] and RAP[29] to generate adversarial examples, we observed a strong intra-family transferability as well as a heterogeneous cross-family transferability which is demonstrated to be irrelevant to involved mechanisms but the distance between models. Here we provide the transferabilty matrices from more methods to confirm our findings. In Fig. 22, we can observe a similar phenomenon as in Fig. 13.
G-B Intra/Cross-training-strategy transferability
In Fig. 23, we expand the set of adversarial example generation candidates for evaluating intra/cross-training-strategy transferability, as discussed in Sec. IV-G, to include NI-FGSM[23], PI-FGSM[27], VT[28], and SI-FGSM[23]. A similar phenomenon is observed: all target models are found more vulnerable to intra-traning-strategy transferability—misclassification of examples generated by surrogate models using the same training strategy.
Figure 27: The rest part of Fig. 26
TABLE IV: Hyper-parameters of decision-based attacks.
@@ G-B Intra/Cross-training-strategy transferability
General hyperparameters maximum perturbation 0.05 0.05 5 1 0.05 0.05 5 1
@@
TABLE V: Hyper-parameters of score-based attacks.
@@
General hyperparameters maximum perturbation 0.05 0.05 5 1 0.05 0.05 5 1
@@
TABLE VI: Hyper-parameters of transfer-based attacks.
@@
$\ell_{\infty}$ $\ell_{2}$ $\ell_{\infty}$ $\ell_{2}$ $\ell_{\infty}$ $\ell_{2}$ $\ell_{\infty}$ $\ell_{2}$bigstrut
General hyper-parameters
maximum perturbation
@@
decay factor / step size of looking ahead
@@
decay factor / step size of looking ahead
@@
late start / iteration number of inner maximization
@@
position to perform linear backpropagation
ResNet-50: the 2nd bottleneck block at the 3rd stage;
${}^{\dagger}$: the conv layer at the 12nd VGG block;
@@
${}^{\dagger}$: the conv layer at the 8th VGG block;
@@
strength of sampled image / number of sampled images
0.2 / 3 0.2 / 3
number of admixed copies / scale factor
@@
intermediate layer to extract feature maps
ResNet-50: the 2nd stage;
${}^{\dagger}$: the conv layer at the 10th VGG block;
@@
${}^{\dagger}$: the conv layer at the 7th VGG block;
@@
ensemble number in aggregate gradient / drop probability
@@
target layer
ResNet-50: the 2nd stage;
${}^{\dagger}$: the conv layer at the 10th VGG block;
@@
${}^{\dagger}$: the conv layer at the 7th VGG block;
@@
intermediate layer
ResNet-50: the 2nd stage;
${}^{\dagger}$: the conv layer at the 10th VGG block;
@@
${}^{\dagger}$: the conv layer at the 7th VGG block;
@@
scale of noise added on DNN weights
ResNet-50: 0.005; ${}^{\dagger}$: 0.01; Inception-V3: 0.01; DenseNet-121: 0.01; ViT-B/16: 0.005;
ResNet-50: 0.005; ${}^{\dagger}$: 0.01; Inception-V3: 0.005; DenseNet-BC: 0.01;
number of ensemble models
@@
parameter of uniform distribution that modulating scalars are drawn from
ResNet-50: 0.22;
@@
batch size / epochs / learning rate / regularization constant
32 / 20 / 0.001 / 6
learning rate schedule
@@
shape-related hyper-parameter of Softplus
ResNet-50: 25
@@
decay factor on residual module
ResNet-50: [1.0, 0.85, 0.65, 0.15]
@@
batch size / epochs / learning rate schedule
512 / 10 / constant 256 / 10 / constant
learning rate/weight decay/momentum
ResNet-50: 0.05/1e-4/0.9; ${}^{\dagger}$: 0.1/7e-4/0.9;
@@
batch size / epochs / learning rate schedule / momentum / $\sigma$
512 / 10 / constant / 0.9 / 0.002 512 / 10 / constant / 0.9 / 0.002
learning rate / weight decay / $\lambda_{\varepsilon,\sigma}$ / $\gamma$
ResNet-50: 0.05 / 5e-4 / 1 / $0.1/\left\|\boldsymbol{\Delta}\mathbf{w}^{*}\right\|_{2}$;
${}^{\dagger}$: 0.005 / 1e-4 / 1 / $0.15/\left\|\boldsymbol{\Delta}\mathbf{w}^{*}\right\|_{2}$/;
@@
${}^{\dagger}$: 0.1 / 1e-4 / 1 / $0.02/\left\|\boldsymbol{\Delta}\mathbf{w}^{*}\right\|_{2}$;
@@
batch size / epochs / learning rate schedule / momentum / $\sigma$
512 / 10 / constant / 0.9 / 0.002 512 / 10 / constant / 0.9 / 0.002
learning rate / weight decay / $\lambda_{\varepsilon,\sigma}$ / $\gamma$
ResNet-50: 0.05 / 5e-4 / 1 / $0.1/\left\|\boldsymbol{\Delta}\mathbf{w}^{*}\right\|_{2}$;
${}^{\dagger}$: 0.005 / 1e-4 / 1 / $0.15/\left\|\boldsymbol{\Delta}\mathbf{w}^{*}\right\|_{2}$/;
@@
${}^{\dagger}$: 0.1 / 1e-4 / 1 / $0.02/\left\|\boldsymbol{\Delta}\mathbf{w}^{*}\right\|_{2}$;
@@
number of sampled models
50 50
rescaling factor of the covariance matrix

### v2

@@ H-A Description of static defenses
H-B Description of dynamic defenses
@@
I Description of analytical tools
In this section, we provide implementation details of our analytical tools.
@@ I Description of analytical tools
FullGrad is a variant of a saliency map that not only captures the importance of each pixel but also provides that of groups of pixels, which correspond to structures. Consider a neural network function $f:\mathbb{R}^{D}\rightarrow\mathbb{R}$ with inputs $\boldsymbol{x}\in\mathbb{R}^{D}$. Let $c_{l}$ be the channel of layer $l$ in a neural network. Let $\psi\left(\cdot\right)$ be the post-processing operations. The FullGrad saliency map is given by
$S_{f}(\boldsymbol{x})=\psi\left(\nabla_{\boldsymbol{x}}f(\boldsymbol{x})\odot\boldsymbol{x}\right)+\sum_{l\in L}\sum_{c\in c_{l}}\psi\left(f^{b}(\boldsymbol{x})_{c}\right),$ (I.1)
where $\nabla_{\boldsymbol{x}}f(\boldsymbol{x})$ is the input-gradient, $f^{b}(\boldsymbol{x})$, short-hand notation for $\nabla_{b}f(\boldsymbol{x},\mathbf{b})\odot\mathbf{b}$, is the bias-gradient. Together, they constitute full-gradients.
@@
$F(\tilde{\boldsymbol{x}})=f(\boldsymbol{x})=f\left(\mathcal{F}^{-1}(\tilde{\boldsymbol{x}})\right).$ (I.2)
@@
$S_{F}(\tilde{\boldsymbol{x}})$ $=\frac{\partial F(\tilde{\boldsymbol{x}})}{\partial\tilde{\boldsymbol{x}}(u,v,c)}$ (I.3)
@@
Adversarial divergence measurement tool When harnessing adversarial transferability from transfer-based attacks, the performance is notably destroyed by the misalignment between a pair of surrogate model $f_{s}$ with parameter $\theta_{s}$ and target models $f_{t}$ with parameter $\theta_{t}$. To properly measure this misalignment, we consider conditional adversarial distribution (CAD), denoted as $\mathcal{P}(\boldsymbol{\delta}_{\boldsymbol{x}}|\boldsymbol{x},\theta)$. As CAD models the distribution of perturbations $\boldsymbol{\delta}_{\boldsymbol{x}}$ conditioned on clean examples $\boldsymbol{x}$ for each model with parameter $\theta$, minimal divergence between CADs of a surrogate-target model pair ensures a high transferability between the surrogate and target models. The rationale behind is that if $\mathcal{P}(\boldsymbol{\delta}_{\boldsymbol{x}}|\boldsymbol{x},\theta_{s})$ and $\mathcal{P}(\boldsymbol{\delta}_{\boldsymbol{x}}|\boldsymbol{x},\theta_{t})$ yield similar probabilities for perturbation $\boldsymbol{\delta}_{\boldsymbol{x}}$, then the perturbation craft by $f_{s}$ can effectively substitute the authentic perturbation crafted by $f_{t}$, thus misleading $f_{t}$. Building on this premise, BlackobxBench quantifies this divergence between $\mathcal{P}(\boldsymbol{\delta}_{\boldsymbol{x}}|\boldsymbol{x},\theta_{t})$ and $\mathcal{P}(\boldsymbol{\delta}_{\boldsymbol{x}}|\boldsymbol{x},\theta_{s})$ as Adversarial Divergence (AD):
$\text{AD}=1-\text{Norm}_{[0,1]}\left(\frac{1}{\log_{2}\hat{D}_{\mathrm{KL}}(s,t)}\right),$ (I.4)
where $\hat{D}_{\mathrm{KL}}(s,t)$ denotes the empirical estimation of KL divergence between $\mathcal{P}(\boldsymbol{\delta}_{\boldsymbol{x}}|\boldsymbol{x},\theta_{t})$ and $\mathcal{P}(\boldsymbol{\delta}_{\boldsymbol{x}}|\boldsymbol{x},\theta_{s})$, $i.e.$, ${D}_{\mathrm{KL}}(s,t)=\left[\mathbb{E}_{\mathcal{P}\left(\boldsymbol{\delta}_{\boldsymbol{x}}\mid\boldsymbol{x},\theta_{s}\right)}\left[\log\frac{\mathcal{P}\left(\boldsymbol{\boldsymbol{\delta}_{\boldsymbol{x}}}\mid\boldsymbol{x},\theta_{s}\right)}{\mathcal{P}\left(\boldsymbol{\boldsymbol{\delta}_{\boldsymbol{x}}}\mid\boldsymbol{x},\theta_{t}\right)}\right]\right]$. $\text{Norm}_{[0,1]}\left(\cdot\right)$ denotes the max normalization. Let $\text{sim}\left(s,t\right)=\frac{1}{\log_{2}\hat{D}_{\mathrm{KL}}(s,t)}$, $\text{Norm}_{[0,1]}\left(\text{sim}\left(s,t\right)\right)=\frac{\text{sim}\left(s,t\right)}{\max_{s,t}\text{sim}\left(s,t\right)}$. AD normalizes the KL divergence to be bounded within $[0,1]$. In the following, we will detail two important parts for computing Adversarial Divergence. We will firstly introduce the modeling of CAD using the conditional generative flow model (c-Glow); then, with the CAD of two models, we will present the computation of KL divergence between them.
Following [41], we learned the c-Glow model for each classifier $f$ to approximate its CAD.
  1. 1.
    Modeling CAD by c-Glow: The c-Glow model learns an invertible mapping between latent variable $\boldsymbol{z}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and perturbation $\boldsymbol{\delta}_{\boldsymbol{x}}$ on benign input $\boldsymbol{x}$:
    $\boldsymbol{\delta}_{\boldsymbol{x}}=g_{\boldsymbol{x},\boldsymbol{\theta}}(\boldsymbol{z}_{0})=g_{\boldsymbol{x},\boldsymbol{\phi}}(\boldsymbol{z}),$ (I.5)
    where $\boldsymbol{\theta}=(\boldsymbol{\phi},\boldsymbol{\mu},\boldsymbol{\sigma})$ is the model parameter. $\boldsymbol{\phi}=\left(\boldsymbol{\phi}_{1},\ldots,\boldsymbol{\phi}_{M}\right)$ is the $M$ layers mapping parameter of c-Glow model which takes $\boldsymbol{z}=\boldsymbol{\mu}+\boldsymbol{\sigma}\odot\boldsymbol{z}_{0}$ as input.
    Then, the conditional likelihood is given by:
    $\log\mathcal{P}_{\boldsymbol{\theta}}(\boldsymbol{\delta}_{\boldsymbol{x}}\mid\boldsymbol{x})$ $=\log\mathcal{P}_{\mathbf{0},\mathbf{1}}\left(\boldsymbol{z}_{0}\right)$ (I.6)
    $+\sum_{i=1}^{M+1}\log\left|\operatorname{det}\left(\frac{\partial g_{\boldsymbol{x},\boldsymbol{\phi}_{i}}^{-1}\left(\boldsymbol{r}_{i-1}\right)}{\partial\boldsymbol{r}_{i-1}}\right)\right|,$
    where $\boldsymbol{r}_{i}=g_{\boldsymbol{x},\boldsymbol{\phi}_{i}}^{-1}\left(\boldsymbol{r}_{i-1}\right)$, $\boldsymbol{r}_{0}=\boldsymbol{\delta}_{\boldsymbol{x}}$, $\boldsymbol{r}_{M}=\boldsymbol{z}$ and $\boldsymbol{r}_{M+1}=\boldsymbol{z}_{0}$.
  2. 2.
    Learning CAD with energy-based model: Directly training c-Glow by maximum likelihood requires collecting a large number of actual adversarial perturbations $\boldsymbol{\delta}_{\boldsymbol{x}}$, which is computationally expensive. To address this, the CAD is instead approximated by an energy-based model:
    $\log\mathcal{P}_{E}(\boldsymbol{\delta}_{\boldsymbol{x}}\mid\boldsymbol{x})\approx-\lambda\cdot\mathcal{L}_{adv}(\boldsymbol{\delta}_{\boldsymbol{x}},\boldsymbol{x}),$ (I.7)
    where $\mathcal{L}_{adv}(\cdot)$ is the adversarial loss function, $\lambda$ is a positive hyper-parameter. Instead of generating adversarial examples, the method samples a large number of perturbations $\boldsymbol{\delta}_{\boldsymbol{x}}$ from the perturbation ball $\mathbb{B}_{\epsilon,\boldsymbol{x}}$ around each clean input $\boldsymbol{x}$. These samples are then scored by the energy model to estimate their importance for learning the CAD.
    The c-Glow model is trained to approximate the energy-based distribution $\mathcal{P}_{E}$ by minimizing the KL divergence:
@@
    $\mathcal{L}=\mathbb{E}_{\mathcal{P}_{E}(\boldsymbol{\delta}_{\boldsymbol{x}}\mid\boldsymbol{x})}\left[\log\frac{\mathcal{\mathcal{P}}_{E}(\boldsymbol{\delta}_{\boldsymbol{x}}\mid\boldsymbol{x})}{\mathcal{\mathcal{P}}_{\boldsymbol{\theta}}(\boldsymbol{\delta}_{\boldsymbol{x}}\mid\boldsymbol{x})}\right].$ (I.8)
The training details can be found in the supplementary materials in [41].
After modeling the CADs of a pair of models, we estimate the KL divergence between them, $i.e.$, $\hat{D}_{\mathrm{KL}}(s,t)$, as follows:
@@
    $\log\frac{\mathcal{P}(\boldsymbol{\delta}_{\boldsymbol{x}}^{(i)}\mid\boldsymbol{x},\theta_{s})}{\mathcal{P}(\boldsymbol{\delta}_{\boldsymbol{x}}^{(i)}\mid\boldsymbol{x},\theta_{t})}.$ (I.9)
@@
Decision boundary visualization tool Decision Surface[103] measures the network’s loss variations concerning the input perturbations and contains the explicit decision boundary, naturally aligning with our analytical requirements of measuring the distance between the benign images and the decision boundary.
@@
$V(i,j,\alpha,\beta)=F(o+i\cdot\alpha+j\cdot\beta).$ (I.11)
For our purpose of assessing the network’s loss variations w.r.t. perturbations, $\alpha$ represents an attack-specific perturbation direction, $\beta$ stands for a random direction vector, and $o$ denotes the benign image. The variables $i$ and $j$ quantify the degree of deviation from point $o$ along the $\alpha$ and $\beta$ directions, respectively.
@@
As shown in the figure of decision surface, with the loss for clean images serving as the central point, the decision surface shows the loss variation w.r.t the perturbation density (shown in x-axis) and the noise density along (shown in y-axis). Evidently, it requires a shorter distance to traverse the decision boundary in the adversarial direction compared to the random direction. This observation indicates that adversarial attacks can identify directions toward the decision boundary within the immediate vicinity of clean images.
@@
TABLE V: Hyper-parameters of decision-based attacks.
@@
General hyperparameters maximum perturbation 0.03 0.03 3 1 0.03 0.03 3 1
@@
TABLE VI: Hyper-parameters of score-based attacks.
@@
General hyperparameters maximum perturbation 0.03 0.03 3 1 0.03 0.03 3 1
@@
TABLE VII: Hyper-parameters of transfer-based attacks.
@@
[…]
position to perform linear backpropagation ResNet-50: the 2nd bottleneck block at the 3rd stage;
VGG-19: the conv layer at the 12nd VGG block;
@@
VGG-19: the conv layer at the 8th VGG block;
@@
strength of sampled image / number of sampled images 0.2 / 3 0.2 / 3
@@
intermediate layer to extract feature maps ResNet-50: the 2nd stage;
VGG-19: the conv layer at the 10th VGG block;
@@
VGG-19: the conv layer at the 7th VGG block;
@@
target layer ResNet-50: the 2nd stage;
VGG-19: the conv layer at the 10th VGG block;
@@
VGG-19: the conv layer at the 7th VGG block;
@@
intermediate layer ResNet-50: the 2nd stage;
VGG-19: the conv layer at the 10th VGG block;
@@
VGG-19: the conv layer at the 7th VGG block;
@@
scale of noise added on DNN weights ResNet-50: 0.005; VGG-19: 0.01; Inception-V3: 0.01; DenseNet-121: 0.01; ViT-B/16: 0.005; ResNet-50: 0.005; VGG-19: 0.01; Inception-V3: 0.005; DenseNet-BC: 0.01;
@@
parameter of uniform distribution that modulating scalars are drawn from ResNet-50: 0.22;
@@
batch size / epochs / learning rate / regularization constant 32 / 20 / 0.001 / 6
@@
shape-related hyper-parameter of Softplus ResNet-50: 25
@@
decay factor on residual module ResNet-50: [1.0, 0.85, 0.65, 0.15]
@@
batch size / epochs / learning rate schedule 512 / 10 / constant 256 / 10 / constant
learning rate/weight decay/momentum ResNet-50: 0.05/1e-4/0.9; VGG-19: 0.1/7e-4/0.9;
@@
batch size / epochs / learning rate schedule / momentum / $\sigma$ 512 / 10 / constant / 0.9 / 0.002 512 / 10 / constant / 0.9 / 0.002
learning rate / weight decay / $\lambda_{\varepsilon,\sigma}$ / $\gamma$ ResNet-50: 0.05 / 5e-4 / 1 / $0.1/\left\|\boldsymbol{\Delta}\mathbf{w}^{*}\right\|_{2}$;
VGG-19: 0.005 / 1e-4 / 1 / $0.15/\left\|\boldsymbol{\Delta}\mathbf{w}^{*}\right\|_{2}$/;
@@
VGG-19: 0.1 / 1e-4 / 1 / $0.02/\left\|\boldsymbol{\Delta}\mathbf{w}^{*}\right\|_{2}$;
@@
batch size / epochs / learning rate schedule / momentum / $\sigma$ 512 / 10 / constant / 0.9 / 0.002 512 / 10 / constant / 0.9 / 0.002
learning rate / weight decay / $\lambda_{\varepsilon,\sigma}$ / $\gamma$ ResNet-50: 0.05 / 5e-4 / 1 / $0.1/\left\|\boldsymbol{\Delta}\mathbf{w}^{*}\right\|_{2}$;
VGG-19: 0.005 / 1e-4 / 1 / $0.15/\left\|\boldsymbol{\Delta}\mathbf{w}^{*}\right\|_{2}$/;
@@
VGG-19: 0.1 / 1e-4 / 1 / $0.02/\left\|\boldsymbol{\Delta}\mathbf{w}^{*}\right\|_{2}$;
@@
number of sampled models 50 50